an intern made a error and gave a pantient a dose of 500mg rahter than 500pg of a drug . what is the percentage of error. must show work
1
Expert's answer
2012-09-13T11:10:51-0400
error mass is 500 pg (petagram)
regular mass is 500 mg
500 pg is
500 000 000 000 000 000 grams and 500 000 000 000 000 000 000 mg

E=I
(Mr-Me)/Mr I = I (500 - 500 000 000 000 000 000 000)/500 I * 100%= 1 x 10^20% of
error